### v2

# FlexLLM: A System for Co-Serving Large Language Model Inference and Parameter-Efficient Finetuning

FlexLLM: A System for Co-Serving Large Language Model Inference and Parameter-Efficient Finetuning
Gabriele Oliaro     Xupeng Miao∗†     Xinhao Cheng     Vineeth Kada     Ruohan Gao     Yingyi Huang
Remi Delacourt     April Yang     Yingcheng Wang     Mengdi Wu     Colin Unger     Zhihao Jia
Carnegie Mellon University      Purdue University      Stanford University
footnotetext: Equal contribution.
Finetuning large language models (LLMs) is essential for task adaptation, yet today’s serving stacks isolate inference and finetuning on separate GPU clusters—wasting resources and under-utilizing hardware. We introduce FlexLLM, the first system to co-serve LLM inference and PEFT-based finetuning on shared GPUs by fusing computation at the token level. FlexLLM ’s static compilation optimizations—dependent parallelization and graph pruning significantly shrink activation memory, leading to end-to-end GPU memory savings by up to 80%. At runtime, a novel token-level finetuning mechanism paired with a hybrid token scheduler dynamically interleaves inference and training tokens within each co-serving iteration, meeting strict latency SLOs while maximizing utilization. In end-to-end benchmarks on LLaMA-3.1-8B, Qwen-2.5-14B, and Qwen-2.5-32B, FlexLLM sustains the inference SLO requirements up to 20 req/s, and improves finetuning throughput by 1.94.8×1.9-4.8$\times$1.9 - 4.8 × under heavy inference workloads and 2.56.8×2.5-6.8$\times$2.5 - 6.8 × under light loads, preserving over 76% of peak finetuning progress even at peak demand. The source code of FlexLLM is publicly available at https://github.com/flexflow/FlexFlow/.
1 Introduction
Recent advancements in large language models (LLMs) such as GPT [12, 24, 57], PaLM [17], and LLaMA [73] have shown strong capabilities of generating natural language texts across various application domains [88, 47, 68, 23]. Given the increasingly high cost of LLM pre-training, there is a trend of standardizing foundational pre-trained LLMs and sharing sub-models across multiple downstream tasks [34, 86] through finetuning (i.e., continuous training on small task-specific datasets). For example, a series of parameter-efficient finetuning (PEFT) techniques [34, 43, 59, 35, 32, 41] have been proposed to update a subset of trainable parameters from an LLM or introduce a small number of new parameters into the LLM while keeping the vast majority of the original LLM parameters frozen. The base LLM sharing paradigm provides opportunities to multiplex resources and improve cluster utilization (e.g., Google [11], Microsoft [35], and Tencent [56]). Many recent efforts [97, 80, 66, 13, 91, 36] inspired by these opportunities to build multi-task serving systems, handling inference requests for multiple fine-tuned LLMs simultaneously in a single GPU cluster.
Another undeniable trend is the increasing importance of LLM finetuning [^1] For simplicity, we use the term finetuning to refer to PEFT for the rest of the paper.. Almost all LLM serving companies, including OpenAI GPT-4o [58], Google Gemini [28], and Databricks Mosaic AI [19], now offer commercial finetuning APIs to facilitate the creation of custom models. This trend is not only due to the widespread adoption of LLMs but also driven by the growing demand for task-specific model customization [49], data access control [69], and the need for continuous and frequent model updating [26]. Unlike inference tasks with strict latency requirements (e.g., microseconds to milliseconds per token), finetuning tasks usually allow for greater latency tolerance with coarser time budget granularity (e.g., minutes to hours). As such, the current practice involves operating these two tasks on dedicated clusters [78, 27, 14], which exclusively use hardware resources through specialized systems optimized for latency and throughput, respectively.
However, real-world LLM inference workloads usually exhibit fluctuating and bursty request arrival patterns [77]. Serving inference requests must meet latency-based service level objectives (SLOs), which may vary from stringent to relaxed depending on the application scenario. To avoid SLO violations, service providers must either over-provision GPU resources or dynamically adjust resource allocation. The former strategy can lead to significant computational waste, while the latter introduces substantial overhead due to the need for frequent redeployments and process switching [10, 54]. In this paper, we explore a key research question: Can we design a system that simultaneously serves both inference and finetuning tasks within a shared GPU cluster, effectively accommodating the time-varying inference computational demands?
We introduce FlexLLM, the first co-serving system for LLM inference and finetuning. FlexLLM introduces a PEFT-as-a-service (PaaS) interface, which unifies inference and finetuning tasks, enabling their joint execution on shared GPU resources. The key insight behind FlexLLM is that inference and finetuning tasks, when using the same base LLMs, can be merged rather than executed separately.
This co-serving approach presents multiple advantages. First, since the base LLM is generally frozen during finetuning, the base model parameters can be stored without redundancy, therefore reducing the memory footprint. Second, when faced with bursty inference demands, FlexLLM can adaptively reassign GPU resources by temporarily reducing the proportion of finetuning tasks, which inherently possess higher latency tolerance. Once the peak inference workload has passed, the finetuning tasks can reclaim resources, thereby optimizing overall throughput and maximizing GPU utilization without compromising inference SLOs. This dynamic strategy provides a novel avenue for cloud LLM service providers to improve operational cost-effectiveness and service capabilities.
Beyond supervised PEFT, a growing class of reinforcement-learning-based fine-tuning algorithms – most notably Group Relative Policy Optimization (GRPO) used in DeepSeek-R1 [31], and multi-objective variants thereof [44] – require the model to generate complete sequences on-the-fly in order to compute policy gradients or reward signals. In each RLHF [65] or GRPO update, the LLM must sample continuations (i.e., perform inference), score them against a learned or hand-crafted reward, and then backpropagate through those same tokens. Because inference and policy-gradient steps are interleaved at the token level, splitting inference and training onto separate GPUs incurs costly data movement and context-switching. This tight coupling further motivates FlexLLM ’s token-level co-serving, which fuses generation and gradient updates in a single pass while still meeting stringent inference SLOs.
Challenges.
Co-serving inference and finetuning tasks within a single system, as envisioned by FlexLLM, presents significant challenges due to their differing workload characteristics. FlexLLM must address three key challenges.
First, how can inference and finetuning tasks be co-located while minimizing the memory footprint? Traditional multi-tenancy ML scheduling systems share resources across tasks and models [81], but suffer from the high memory demands associated with LLMs. Furthermore, existing finetuning systems usually retain all intermediate activations similar to training systems to enable gradient computation during backpropagation, which substantially increases their resource consumption compared to pure inference tasks. This discrepancy makes it challenging to manage system resources between finetuning and inference tasks.
To address this challenge, FlexLLM co-locates finetuning tasks with inference tasks by enforcing them to share the base LLM parameters and intermediate activation buffers. In scenarios where the base LLM has already been deployed in a distributed system, FlexLLM uses dependent parallelization to optimize the placement of small, trainable bypass networks (e.g., LoRA adapters [35]) introduced during finetuning. This is achieved by solving a constrained optimization problem detailed in §5.1.
Unlike existing systems that preserve all intermediate activations throughout a forward pass for finetuning, FlexLLM leverages the observation that most intermediate activations are only necessary for computing gradients of frozen parameters. By implementing graph pruning, FlexLLM eliminates unnecessary data dependencies associated with these frozen parameters (see §$\S$ 5.2). These memory optimization strategies allow FlexLLM to reduce the memory overhead associated for finetuning by up to 8$\times$ compared to existing systems.
Next, how can we run inference and finetuning tasks simultaneously while minimizing the interference with inference latency and SLO attainment? Previous work on traditional ML workloads [62] has shown that co-locating inference with training can significantly degrade inference performance. Unlike conventional training workloads that require coarse-grained temporal or spatial resource sharing [15], LLM finetuning provides a unique opportunity to batch base LLM computation and parameter access across requests from different finetuned variants [80]. However, due to differences in execution logic—auto-regresive decoding in inference versus backpropagation in finetuning—merging the two types of tasks remains difficult.
To overcome this challenge, FlexLLM introduces a novel token-level finetuning mechanism, which decomposes the finetuning computation of a sequence into smaller, token-level units (§$\S$6.1). During the forward pass, finetuning tokens follow the same operator execution order as decoding tokens in inference requests. These tokens are then processed using fused GPU kernels. For the backward pass, FlexLLM launches separate GPU streams for finetuning tokens and adopts a layer-wise execution strategy to minimize memory usage. This design significantly reduces the performance impact of finetuning tokens on inference latency, enabling efficient co-serving of inference and finetuning tasks on shared hardware.
Finally, how can we handle fluctuating inference workloads while preserving their latency SLO? Given the dynamic nature of inference request arrival patterns, provisioning for peak demand often results in significant resource under-utilization. FlexLLM’s fine-grained, token-level execution strategy enables quick adaptation to load fluctuations. Inspired by prior work on co-locating multiple cloud workloads [25], FlexLLM prioritizes inference tokens to meet strict latency constraints, while inserting finetunign tokens in a best-effort manner.
To achieve this, FlexLLM introduces a hybrid token scheduler, which coordinates the scheduling of inference and finetuning tokens at each co-serving iteration. This scheduler aims to maximize overall GPU utilization while guaranteeing that inference tasks consistently meet their SLOs (Section 6.2).
Evaluation.
We compare FlexLLM’s co-serving approach with existing approaches that employ separate GPUs to perform inference and finetuning and show that FlexLLM achieves significantly higher finetuning throughput. For example, under heavy inference workloads where existing systems must dedicate all available GPUs for inference tasks and therefore cannot make any finetuning progress, FlexLLM can still preserve more than 80% of peak finetuning throughput.
2 Background and Challenges
2.1 Parameter-Efficient Finetuning
Finetuning enables LLMs to adapt to specific domains and downstream tasks [20, 61, 12]. However, full finetuning of an LLM introduces significant computational and memory overheads due to the model’s large parameter count. To address this issue, a class of parameter-efficient finetuning (PEFT) methods [22] has been recently proposed. Inspired by the sensitivity of LLMs to input prompts [63], a series of studies introduce trainable prompt embeddings—either prepended to the input sequence or incorporated into attention components—while keeping the original LLM parameters frozen [48, 43, 41]. Recent work also proposed a variety of adapter modules to insert lightweight task-specific modules into the Transformer layers. For example, LoRA [35] leverages the low-rank decomposition concept to construct trainable parameters inserted into the original LLM weights. (IA)3 [46] proposes to scale the pre-trained model weights of an LLM with a trainable vector. Recently, there has been a surge in the proposal of unified approaches that amalgamate various PEFT methods by leveraging human heuristics [32] or employing neural architecture search [96, 98, 51]. Existing PEFT approaches focus on optimizing model performance while minimizing trainable parameters. However, reducing the number of trainable parameters does not inherently imply a corresponding reduction in memory footprint.
Figure 1: Comparing different resource sharing approaches for serving finetuning and inference. For spatial sharing and co-serving, the height of rounded rectangles illustrates the splitting ratio of GPU resources (e.g., streaming multi-processors).
2.2 PEFT-based LLM Inference
LLM inference is increasingly essential in both research and industrial applications. Recent work has introduced a diversity of systems for simultaneously serving multiple PEFT-based LLMs. PetS [97] is the first unified framework for multi-task parameter-efficient Transformers serving, which decouples linear layers in the computation graph into task-agnostic shared operations and task-specific PET operations, then optimizes the operator scheduling strategy to improve concurrency and throughput. While PetS only supports early encoder-only models, recent work also focuses on optimizations for generative decoding models. Punica [13] is a system to serve multi-tenant LoRA in a shared GPU cluster. It proposes a new CUDA kernel called Segmented Gather Matrix-Vector Multiplication, which allows batching GPU operations for the concurrent execution of multiple, different LoRA models. S-LoRA [66] is optimized on the kernel introduced by Punica. It proposes unified paging, heterogeneous batching, and a novel tensor parallel algorithm to serve thousands of LoRA components on a single machine. dLoRA [80] further improves efficiency by dynamically adjusting the LoRA batching plans under imbalanced workloads.
2.3 Reinforcement-Learning Based Finetuning
While PEFT methods such as LoRA and adapters optimize a small set of task parameters via supervised losses, recent work has demonstrated that pure reinforcement learning (RL) can substantially enhance LLM reasoning and alignment. In RLHF (e.g., [67]), the model samples candidate completions, a human or reward model ranks them, and policy gradients adjust the generator to favor preferred outputs. More recently, Group Relative Policy Optimization (GRPO) forgoes a separate value network by scoring groups of sampled continuations to estimate baselines, reducing critic overhead [31]. Extensions such as multi-objective GRPO further handle safety and helpfulness rewards within a single RL loop [44], and open-source systems like DAPO illustrate scaling these methods to 32B-parameter models [85].
Crucially, all of these RL-based pipelines require on-the-fly LLM generations as part of the reward calculation, interleaving inference kernels with backward passes at sub-sequence granularity. This pattern conflicts with the traditional separation of “inference-only” versus “training-only” clusters, leading to GPU underutilization or high-latency context switches. FlexLLM ’s token-level fusion directly addresses this challenge by co-scheduling generation and gradient updates within a single GPU pipeline, while still meeting strict inference SLOs.
3 Co-serving LLM Inference and PEFT
Previous studies have proposed different methods for scheduling heterogeneous workloads (in our case, inference and PEFT) on shared GPU resources. These methods can be classified into three categories: resource isolation, temporal sharing, and spatial sharing. This section discusses the limitations inherent to these conventional approaches when applied to co-serving LLM inference and PEFT. It also highlights the fundamental distinctions between FlexLLM’s co-serving solution and previous work. Figure 1 depicts the GPU execution timelines associated with different scheduling approaches for managing inference and finetuning tasks.
Resource isolation.
A common approach to support both inference and finetuning tasks is resource isolation, which partitions the available GPUs into two distinct groups, each dedicated to handling one type of workload (see Figure 1(a)). This approach ensures that throughput-intensive finetuning tasks do not interfere with latency-critical inference tasks. When finetuning an LLM, a dataset of requests is provided to train the PEFT layers, with all requests submitted to FlexLLM simultaneously, making it easier to build larger batches that can be processed sequentially. In contrast, inference requests are submitted independently by users who expect timely processing within predetermined SLOs, with the arrival times of these requests unpredictable.
Resource isolation can introduce significant GPU under-utilization, particularly in scenarios with variable request arrival rates. This is because finetuning and inference requests cannot be batched together if the number of pending requests of either type is insufficient to fully utilize a batch. With fewer GPUs available to parallelize inference computations, it may not be possible to serve requests fast enough to meet their SLO (e.g., request $r_{3}$ in Figure 1(a)).
Temporal sharing.
Compared to resource isolation, temporal sharing [81, 82] improves GPU utilization by allowing different tasks to share GPU resources through time slices of different lengths depending on various parameters. This method allows both inference and finetuning requests to access all GPUs within the cluster in turns, as shown in Figure 1(b). By using all GPUs to parallelize computation (e.g., with data parallelism or tensor model parallelism), the latency for each request can be significantly reduced, thus aiding in meeting inference SLOs. For example, while inference request $r_{3}$ fails to complete in the scenario illustrated in Figure 1(a) due to resource limitations, it successfully executes in Figure 1(b) by leveraging more resources. However, the arrangement of time slices can also be harmful because of the unpredictable arrival times of inference requests, which may have a stringent latency requirement that cannot be met (e.g., $r_{4}$ has been delayed in Figure 1(b) by pre-scheduled finetuning requests).
Spatial sharing.
Spatial sharing offers an alternative by enabling simultaneous execution of different tasks at the kernel level, where each task utilizes a specific portion of the GPU resources (i.e., streaming multi-processors) This can be implemented on NVIDIA GPUs through multi-stream programming or by utilizing the Multi-Process Service (MPS) [4]. Furthermore, on the latest Hopper and Ampere architectures, spatial sharing benefits from the Multi-Instance GPU (MIG) [3], which allows for dynamic adjustments of GPU partition configurations. However, these adjustments often lack the flexibility and efficiency required to handle dynamic inference workloads effectively [21, 83, 2]. Figures 1(c) and (d) show two scenarios where a substantial portion of GPU resources is allocated to either finetuning or inference tasks. While over-provisioning for inference tasks may appear to provide better SLO guarantees, it still falls short during peak demand periods such as the unexpected urgent request $r_{3}$. Additionally, over-provisioning wastes resources and significantly slows down the finetuning process. Therefore, relying solely on spatial sharing to manage finetuning and inference tasks often results in suboptimal performance  [15, 92].
FlexLLM’s co-serving approach.
As shown in Figure 1(e), FlexLLM employs a co-serving approach that uses a fine-grained scheduling mechanism to adaptively manage both inference and finetuning requests. During each co-serving iteration, FlexLLM jointly serves inference and finetuning requests by (1) computing keys and values for inference requests in the pre-filling phase, (2) generating a new token for inference requests in the decoding phase, and (3) performing forward and/or backward passes for selected tokens from finetuning requests.
Unlike temporal or spatial sharing approaches, which schedule GPU kernels of existing systems independently, FlexLLM integrates inference and finetuning kernels. This integration reduces the overhead associated with kernel launches and minimizes accesses to GPU device memory to retrive model weights. In each iteration, FlexLLM dynamically adjusts the allocation of inference and finetuning tokens, strategically balancing the need to meet the SLOs of inference requests with the goal of maximizing GPU utilization. This adaptive approach ensures that FlexLLM adheres to the requirements of inference requests while optimizing the use of available GPU resources.
4 System Overview
Figure 2: An overview of FlexLLM.
Figure 2 shows an overview of FlexLLM, which co-serves both PEFT and inference requests for LLMs. Users can choose to use either finetuning or inference services on a model from the PEFT model hub, which stores the backbone LLM and all finetuned variants. The programming interface for both inference and finetuning requests is unified through a PEFT-as-a-Service (PaaS) interface. FlexLLM comprises three main components. First, for finetuning requests that execute in parallel across multiple GPUs, the static compilation module (§5) generates a parallel computation graph. This graph specifies the execution of the PEFT model over the distributed environment and optimizes the graph by pruning unnecessary tensors for memory saving. Second, the dynamic scheduling module (§6) adapts a token-level finetuning mechanism. It mixes inference and finetuning tokens with a hybrid scheduling policy. Finally, the computation graph and schedule plan are executed by FlexLLM’s distributed co-serving runtime (§7).
Figure 3: Four possible parallel states for a tensor dimension and their transitions. For each parallel state, the symbol in parenthesis shows the notation FlexLLM used to represent it.
Figure 4: Illustration of FlexLLM’s different dependent parallelization strategies with the LoRA example. Each green (or gray) box indicates a compute (or parallelization) operator, and each edge between operators represents a parallel tensor, and the parallelization states of the tensor’s dimensions are shown next to the edge.
4.1 PEFT-as-a-Service Interface
To facilitate the sharing of the backbone LLM across different tasks, FlexLLM represents a PEFT model as a sequence of bypass networks attached to the backbone LLM. Each bypass network takes a single tensor from the backbone LLM as input and produces a single output tensor, which is added to one tensor of the backbone LLM. Let $X$ and $Y$ denote the input and output tensor of the bypass network, and let $f_{B}(X)$ and $f_{A}(X)$ denote the neural architecture of the backbone and bypass network for calculating $Y$. The bypass network can be formulated as $Y=f_{B}(X)+f_{A}(X)$. All existing PEFT methods can be represented in this format. For instance, the (IA)3 architecture modifies the topology of the backbone LLM by inserting an elementwise multiplication operator. It can be transformed into FlexLLM’s bypass network format since $Y=X\odot W=X+X\odot(W-O)$, where $\odot$ is elementwise multiplication and $O$ is a matrix of the same shape as $W$ whose elements are all one. A significant benefit of this interface is that all PEFT models preserve the neural architecture of the backbone LLM, which enables FlexLLM to fuse the computation graphs of different PEFT models. FlexLLM provides a unified user interface to launch inference and finetuning tasks by specifying a PEFT model and committing different types of input data, such as finetuning datasets or inference prompts.
5 Static Compilation
Once a PEFT model is registered, FlexLLM compiles it into a parallel computation graph (PCG) [75]. To optimize finetuning performance, FlexLLM applies two key static compilation optimizations to discover an optimized PCG: dependent parallelization5.1) and graph pruning5.2).
5.1 Dependent Parallelization
To minimize memory overhead, FlexLLM allows finetuning and inference tasks to utilize the same backbone LLM and share memory allocated for both model weights and intermediate activations. This approach avoids redundant memory consumption while requiring the parallelization strategies of the PEFT models to be compatible with that of the backbone LLM, a task termed dependent parallelization.
Existing auto-parallelization approaches like Alpa [93] and Unity [75] can identify optimal parallelization strategies for backbone LLMs. FlexLLM extends the optimization target to also include the bypass networks in PEFT models. To achieve this goal, FlexLLM generalizes the parallel computation graph (PCG) abstraction introduced in Unity [75] by associating a state field with tensor dimension.
Figure 3 shows the four possible tensor states when parallelizing a tensor, including non-parallel (-), partitioned (|), replicated (=), and pre-reduce (+), as well as the state transitions. Given the fixed parallelization of the backbone LLM, FlexLLM searches for an optimal PCG for each bypass network by enumerating possible parallelization operators—partition, combine, replicate, and reduce—inferring output tensor states, and validating their compatibility. Figure 4 shows four candidate PCGs discovered by FlexLLM for parallelizing a LoRA network. To select the best strategy, FlexLLM reuses Unity’s profiling-based cost model [75] and chooses the candidate PCG with the lowest estimated execution cost.
Algorithm 1 Static graph pruning. For an operator $n$, $\textsc{UpdateInput}(n,\mathcal{O}(n))$ returns a set of input tensors needed in order only to compute $\mathcal{O}(n)$ of the operator.
1:Inputs: PCG of a PEFT model $\mathcal{G}$
2:Outputs: $\mathcal{A}$ is a set of tensors to be memorized, $\mathcal{R}$ is a set of tensors to be rematerialized
3$\triangleright$ Step 1: computation graph pruning
[…]
5:$\mathcal{Q}=\emptyset$triangleright $\mathcal{Q}$ is a queue of updated operators
6:for operator $n\in\overline{\mathcal{G}}$ do
7:     for output tensor $t\in\mathcal{O}(n)$ do
8:         if $t$ is the weight gradient of the base LLM then
9:              $\mathcal{O}(n)=\mathcal{O}(n)\setminus\{t\}$
10:              $\mathcal{I}(n)=\textsc{UpdateInput}(n,\mathcal{O}(n))$
11:              $\mathcal{Q}$.push_back($n$)
12:while $\mathcal{Q}$ is not empty do
13:     $n=\mathcal{Q}$.pop_front()
14:     for output tensor $t\in\mathcal{O}(n)$ do
15:         if $\nexists u.t\in\mathcal{I}(u)$  then
16:              $\mathcal{O}(n)=\mathcal{O}(n)\setminus\{t\}$
17:              $\mathcal{I}(n)=\textsc{UpdateInput}(n,\mathcal{O}(n))$
18:              $\mathcal{Q}$.push_back($n$)
19:$\mathcal{A}=\emptyset$
20:for operator $n\in\mathcal{G}$ do
21:     for tensor $t\ in\mathcal{O}(n)$ do
22:         if $\exists u\in\overline{\mathcal{G}}.t\in\mathcal{I}(u)$ then
23:              $\mathcal{A}=\mathcal{A}\cup\{t\}$
24$\triangleright$ Step 2: opportunistically rematerializing tensors
25:for tensor $t\in\mathcal{A}$ do
26:     Let $n$ be the operator that outputs $t$ (i.e., $t\in\mathcal{O}(n)$)
27:     if $\mathcal{I}(n)\subseteq\mathcal{A}$ and $\textsc{Cost}(n)<threshold$ then
28:         $\mathcal{A}=\mathcal{A}\setminus\{t\}$, =$\mathcal{R}=\mathcal{R}\cup\{t\}$
29:return $\mathcal{A},\mathcal{R}$
Figure 5: Static graph pruning for an MLP model with LoRA.
Figure 6: An overview of existing parameter-efficient fine-tuning (PEFT) methods. Green boxes show the operators of the backbone LLM, while blue boxes are the operators introduced by different PEFT methods. Arrows represent intermediate activations to be reserved and the dashed arrows demonstrate they are pruned by FlexLLM.
5.2 Graph Pruning
Figure 7: Illustration of attention module’s forward and backward execution with FlexLLM’s token-level finetuning mechanism.
FlexLLM’s graph pruning algorithm takes as input a PEFT model and its backbone LLM, and outputs a minimal set of intermediate activations that must be reserved to perform finetuning of the bypass networks. This pruning process requires reasoning over data dependencies between operators, specifically their input and output tensors. To facilitate this analysis, FlexLLM introduces additional notation in the PCG representation. Let $\mathcal{G}=(\mathcal{N},\mathcal{E})$ denote the PCG of a PEFT model, where each node $n\in\mathcal{N}$ represents a tensor algebra (or parallelization) operator, and each edge $e=(n_{1},n_{2})\in\mathcal{E}$ is a tensor shared between operators. For a node $n$, let $\mathcal{I}(n)$ and $\mathcal{O}(n)$ denote its set of input and output tensors. By definition, $(n_{1},n_{2})\in\mathcal{E}$ if and only if $\mathcal{O}(n_{1})\cap\mathcal{I}(n_{2})\neq\emptyset$, i.e., when at least one output tensor of $n_{1}$ serves as an input to $n_{2}$.
The design of FlexLLM’s graph pruning algorithm is based on two key observations. First, conventional ML training procedures maintain all intermediate activations during the forward pass to compute gradients in the backward pass. However, due to the linear algebra nature of most operators, these intermediate activations are primarily used to compute gradients with respect to the base LLM’s trainable parameters. In the case of PEFT, the bacebone LLM parameters are frozen, and only the parameters in the bypass networks are updated. This insight allows FlexLLM to eliminate most intermediate activations, retaining only those necessary for computing gradients of the bypass networks. Second, different PEFT methods connect bypass networks at varying locations within the backbone LLMs, and therefore require different subsets of intermediate activations for backpropagation. Figure 6 illustrates examples of how various PEFT methods attach their bypass networks to the base LLM.
Building on these observations, FlexLLM performs static graph pruning during the construction of the PCG for each registered PEFT model, and dynamical scheduling to manage finetuning and inference tasks, along with corresponding memory allocations (see Section 6). Algorithm 1 shows FlexLLM’s graph pruning algorithm. Given a PEFT model’s PCG $\mathcal{G}$, the algorithm generates an execution plan that specifies whether each intermediate activation should be cached during the forward pass to enable efficient backpropagation.
First, ReverseAutoDiff constructs the backward graph $\overline{\mathcal{G}}$ using reverse-mode automatic differentiation [33]. Since all weights of the base LLM are frozen during finetuning, FlexLLM prunes gradients the computation of their weights. Pruning the output tensors of an operator may, in turn, create opportunities to prune downstream operators. FlexLLM maintains a queue $\mathcal{Q}$ of operators whose input and/or output tensors have changed. It iteratively processes this queue to discover further pruning opportunities. Upon completion, FlexLLM identifies a set of tensors $\mathcal{A}$ representing the minimal set of intermediate activations that must be preserved for the backward pass. FlexLLM further combines graph pruning with two established memory optimizations: rematerialization and activation compression.
Rematerialization.
Tensor rematerialization [37] consists of selectively discarding some tensors in the forward pass and recomputing them during backward propagation. Originally designed to reduce the peak memory requirement of DNN training, this technique is also beneficial for LLM co-serving. For each remaining tensor $t\in\mathcal{A}$ after graph pruning, FlexLLM rematerializes $t$ if and only if (1) all input tensors of $t$ are stored in $\mathcal{A}$ and (2) recomputing $t$ using the input tensors introduces a low rematerialization overhead.
Activation compression.
In addition, FlexLLM also opportunistically applies lossless compression for remaining tensors whenever possible. This is because performing backward computation for some operators does not require access to the original input tensors. One example is the ReLU operator $y=\texttt{ReLU}(x)=\max{(x,0)}$ whose derivative is $\partial y/\partial x=1$ for $x>0$ and 00 for $x\leq 0$. Therefore, instead of storing the original input tensor $x$, FlexLLM keeps the bitmask of $x$.
6 Dynamic Scheduling
Once static compilation produces an executable generalized PCG for a PEFT model, FlexLLM must determine an execution schedule capable of concurrently handling both finetuning and inference requests. This is particularly challenging due to the distinct characteristics of these two workloads. First, inference requests arrive in an online and highly dynamic manner, requiring the system to rapidly adapt GPU resource allocation to maintain low-latency responses. In constrast, finetuning involves both a forward pass—similar to inference—and a backward pass to update model parameters. Additionally, auto-regressive inference for generative LLMs proceeds in a token-by-token fashion, while finetuning is typically performed at the sequence level to maximize throughput. To address these challenges, FlexLLM introduces a novel token-level finetuning mechanism along with a hybrid token scheduler.
Figure 8: Illustration of KV gradients accumulation in the backward pass of FlexLLM’s token-level finetuning mechanism, where $s_{j}$ and $l_{j}$ are the window length and the starting position of the $j$-th slice of the sequence.
6.1 Token-Level Finetuning
Processing a fine-tuning request typically involves performing both forward and backward passes across all model layers for the entire input sequence. When co-served with inference requests, this sequence-level execution can significantly interfere with the token-wise inference kernels and degrade system goodput—this is, the effective inference throughput that meets latency SLOs.
To address this problem and maximize GPU utilization, FlexLLM introduces a token-level finetuning mechanism, detailed in Algorithm 2. The key idea is to decompose the finetuning computation into smaller steps using a dynamic sliding window over tokens. This window spans both the forward and backward passes, with its size dynamically determined by the hybrid token scheduler (line 4 and 15) to ensure that inference requests continue to meet their SLO constraints (§6.2).
During the forward pass (lines 3-11), the finetuning sequence is partitioned into windows of $s_{i}$ tokens (line 5), which are incrementally fed into the model. The model processes each token window layer by layer (lines 7-8), computes the generative loss for the window (line 10), and then advances to the next set of tokens (line 11). To preserve the semantic of full sequence-level finetuning, FlexLLM caches key and value tensors—similar to incremental decoding in inference [84]—as well as query tensors, which are reused during backward attention computations. This caching follows the causal mask constraint inherent in LLMs.
During the backward pass (lines 12-21), FlexLLM executes the model layers in reverse order (line 13). Within each layer, the finetuning request is again divided into token windows (lines 14-16). However, unlike the relatively straightforward forward pass, token-level backward execution of the attention module is more complex due to auto-regressive token dependencies.
We illustrate the forward and backward PCGs of the attention module in Figure 7, with the forward graph on the left and the backward graph on the right. When computing gradients for a slice of $s_{j}$ tokens, the resulting gradients for the queries, keys, and values—denoted as $\Delta Q$, $\Delta K$, $\Delta V$)—have shapes of $[s_{j},h]$, where $h$ is the hidden size. To compute $\Delta Q$, the model must access the cached keys from all preceding $l_{j}$ tokens, consistent with the causal mask in forward pass.
However, due to the autoregressive nature of attention, the loss from the $s_{j}$ tokens also contributes to gradients with respect to previous tokens. As a result, both $\Delta K$ and $\Delta V$ have shapes of $[l_{j},h]$, encompassing all tokens prior to and including the current window. As shown in Figure 8, FlexLLM accumulates these KV gradients across steps. Only after the full backward traversal of the sequence is complete does FlexLLM apply the accumulated gradients (of size $[s_{j},h]$) to compute gradients for subsequent model parameters. Note that, this accumulation strategy only slightly increases memory consumption because FlexLLM’s layer-wise execution design allows each layer to reuse the workspace memory allocated by the previous layer, ensuring efficient memory utilization during backward pass.
Algorithm 2 Token-level finetuning mechanism.
1:Inputs: PEFT model $M$ depth of $N$; Fine-tuning request $r$ with sequence length of $L$
2:Outputs: PEFT model gradients $\Delta M$
3:for $i\leftarrow 0,l_{0}\leftarrow 0$; $l_{i}<L$; $i\texttt{++}$+ do $\triangleright$ Forward pass
4:     $s_{i}\leftarrow\textsc{HybridTokenScheduler}(l_{i},L)$
5:     $r_{i}\leftarrow\textsc{Slice}(r,l_{i},s_{i})$
6:     $X_{0,i}\leftarrow\textsc{Tokenize}(r_{i})$
7:     for model layer index $n\in\textsc{Range}(0,N)$ do
8:         $X_{n+1,i},Q_{n,i},K_{n,i},V_{n,i}\leftarrow\textsc{Forward}(M_{n},X_{n,i})$
9:         $QKVCache_{n}\leftarrow\textsc{Append}(Q_{n,i},K_{n,i},V_{n,i})$
10:     $Loss_{i}\leftarrow\textsc{GenerativeLoss}(X_{N,i},r_{i})$
11:     $l_{i+1}\leftarrow l_{i}+s_{i}$
12:$Y_{N}\leftarrow Loss$
13:for model layer index $n\in\textsc{Range}(N-1,-1,-1)$ do
14:     for $j\leftarrow 0,l_{0}\leftarrow L$; $l_{j}>0$; $j\texttt{++}$+ do $\triangleright$ Backward pass
15:         $s_{j}\leftarrow\textsc{HybridTokenScheduler}(l_{j},0)$
16:         $Y_{n+1,j}\leftarrow\textsc{Slice}(Y_{n+1},l_{j},s_{j})$
17:         $Y_{n,j},\Delta K_{n,j},\Delta V_{n,j}\leftarrow\textsc{Backward}(M_{n},Y_{n+1,j},$
[…]
19:         $\Delta KVAccum_{n}\leftarrow\textsc{Add}(\Delta K_{n,j},\Delta V_{n,j})$
20:         $G_{n,j}\leftarrow\textsc{Slice}(\Delta KVAccum_{n},l_{j},s_{j})$
21:         $\Delta M_{n}\leftarrow\Delta M_{n}+\textsc{CalculateGrads}(M_{n},G_{n,j})$
22:     $l_{j+1}\leftarrow l_{j}-s_{j}$
23:return $\Delta M$
Figure 9: Execution timeline and token scheduling when co-serving inference requests and a finetuning mini-batch.
6.2 Hybrid Token Scheduler
FlexLLM’s hybrid token scheduler coordinates token-level GPU execution in two stages. First, FlexLLM determines the scheduling of available inference requests based on specific scheduling policies. By default, FlexLLM adopts Orca’s iteration-level scheduling [84], which maintains a fixed maximum batch size and dynamically replaces each completed request with a new one whenever available. To further mitigate blocking caused by long input sequences, FlexLLM incorporates the chunked-prefill optimization [6].
Second, after determining the inference schedule for a given iteration, FlexLLM opportunistically appends as many finetuning tokens as possible—determined by the sliding window size $s$—to maximize GPU utilization. The number of finetuning tokens added is determined automatically using the formula $s=\operatornamewithlimits{arg\,max}f(c,s)\leq$SLO, where $f(\cdot,\cdot)$ is the latency estimation function and $c$ is the number of inference tokens scheduled in the current iteration. Here $f(\cdot,\cdot)$ is derived via offline profiling of the LLM’s execution [55]. Such window-based, best-effort scheduling mechanism ensures that FlexLLM can automatically adapt to inference workload fluctuations, maintaining inference SLOs while opportunistically allocating compute to finetuning tasks (§8.3).
Figure 9 illustrates the execution timeline of FlexLLM and shows an example a token scheduling plan for a finetuning mini-batch $A$, under a representative inference requests arrival pattern (i.e., $r_{0}$ to $r_{4}$). To preserve the semantics of finetuning, FlexLLM enforces the execution dependencies between the forward and backward passes of each mini-batch. During the forward pass, FlexLLM leverages fused GPU kernels to jointly process both inference and finetuning tokens. This approach avoids additional kernel launch overhead and is enabled by the shared token-wise computation logic between inference and finetuning. In particular, finetuning tokens follow the same execution pattern and causal masking rules as inference tokens during the prefill phase, guaranteeing compatibility at the kernel level. For the backward pass, FlexLLM adapts a space-sharing approach, using two separate GPU streams: one for inference tokens and another for finetuning tokens. This design allows concurrent execution without interference.
Figure 10: End-to-end comparison between co-serving and using separate resources on three models.
Figure 11: End-to-end comparison between co-serving
7 Implementation
FlexLLM is built on top of FlexFlow Serve [53] with 9K lines of C++ and CUDA code. It comes with both C++ and Python bindings, and its Python interface is fully compatible with the HuggingFace’s PEFT [50] and transformers [79] libraries. This means that users can easily serve models imported directly from HuggingFace, and even upload finetuned PEFTs to the HuggingFace Hub. Moreover, FlexLLM supports any PEFT models with LLAMA [73], GPT [12], OPT [90], Falcon [9], or MPT [70] backbones. In addition, we provide instructions on how to add support for custom models from HuggingFace.
Memory management.
FlexLLM manages the required GPU memory using a combination of static and dynamic allocation. Static allocation reserves GPU space for the backbone weights and the KV cache required for incremental decoding during inference. Dynamic allocation is used for gradients, activations, and optimizer states. When a finetuning request is received, FlexLLM allocates memory during the first forward pass and reuses it during subsequent backward passes, freeing up space when it’s no longer needed.
Key-value gradient accumulator.
To support token-level finetuning, it is necessary to accumulate key-value gradients for every preceding token in a given sequence during backpropagation. FlexLLM utilizes static allocation to reserve space for key-value gradient accumulation. During backpropagation, the gradients for keys and values of a specific segment are obtained, and the position of this segment in the overall sequence is determined. FlexLLM then accumulates the new key-value gradients of all subsequently scheduled segments. When a segment is scheduled during backpropagation, it includes accumulated gradients from future tokens.
8 Evaluation
LLMs and PEFT models.
We evaluated the performance of FlexLLM by comparing it with existing approaches across three publicly available LLMs: LLaMA-3.1-8B [74], Qwen-2.5-14B [60], and Qwen-2.5-32B [60]. Pre-trained model weights were obtained from their respective HuggingFace repositories [1]. Since existing LLM serving systems only support base models and LoRA-based PEFT, our evaluation focuses on these two configurations to ensure fair comparison. Following prior work, we applied LoRA with a rank of 16 to the down projection layers in each multi-layer perceptron (MLP) block of the models. This setup yields 9.4M, 14.5M, and 25.16M trainable parameters for the three LLMs.
Platform.
[…]
Workload.
To synthesize the inference workloads, we sampled the prompt lengths and generation lengths of each request from the ShareGPT [72] dataset, following prior work [39, 95, 7]. However, ShareGPT does not include request arrival times. To address this limitation, we used arrival time data from a recent production trace [76] collected from ChatGPT serving on Azure. The original trace spans multiple days. To evaluate the online serving capabilities of FlexLLM, we sampled 20-minute intervals from the trace and replayed the requests, adjusting the arrival times to simulate different average arrival rates. We used a per-token inference latency of 50ms, 75ms, and 75ms as the SLO for inference requests, following prior work [7]. We set the time-to-first-token (TTFT) SLO to 5s per request to prevent excessive queueing.
We sampled each finetuning request from the NovaSky-AI reasoning dataset [71], and truncate each sequence length to a maximum of up to 8192 tokens. We used the Adam optimizer for training [38].
8.1 End-to-end Comparison
In this section, we compare FlexLLM’s co-serving approach with existing approaches that employ separate clusters for inference and finetuning tasks. In the separate cluster approach, we use vLLM [39] and LlamaFactory [94] as the inference and finetuning systems, respectively, as they represent state-of-the-art solutions for these workloads. In this experiment, we allocated 4, 8, and 16 A100 GPUs for LLaMA-3.1-8B, Qwen-2.5-14B and Qwen-2.5-32B, respectively, to serve inference and finetuning workloads.
Figure 10 shows the results. The separate cluster approach applied a tensor parallel degree of 1, 2, and 4 for the three models, respectively, creating an overall of four inference and finetuning pipelines. We explored various strategies to distribute resources between inference and finetuning tasks. For example, a configuration of 25% vLLM - 75% LlamaFactory indicates that 25% of the resources (i.e., one pipeline) are dedicated to inference, while 75% (i.e., 3 pipelines) are for finetuning.
Across all three models, configurations using one or two pipelines for inference—25% vLLM and 50% vLLM—are only capable of managing lightweight inference workloads, achieving limited inference throughput and SLO attainment under more intensive workloads. Conversely, dedicating three pipelines to inference (75% vLLM - 25% LlamaFactory) maintains high SLO attainment for inference tasks but results in suboptimal finetuning throughput.
FlexLLM’s co-serving approach maintains near-optimal SLO attainment but without compromising finetuning performance. By prioritizing latency-sensitive inference tasks in each co-serving iteration and opportunistically appending as many finetuning tokens as possible, FlexLLM maximizes GPU utilization and improves finetuning throughput.
Across all three models, FlexLLM matches the 75% vLLM - 25% LlamaFactory configuration in inference SLO attainment (at or above 90% even at 20 req/s) and inference throughput, while dramatically improving finetuning throughput. Specifically, under heavy inference loads (20 req/s), FlexLLM sustains finetuning throughputs of 7.2K, 2.2K and 2.2K tokens/s for LLaMA-3.1-8B, Qwen-2.5-14B, and Qwen-2.5-32B respectively, compared to only 3.8K, 1.0K, and 0.5K tokens/s in the 75% vLLM - 25% LlamaFactory setup—i.e. a 1.9$\times$-4.8$\times$ improvement. Under light inference loads (4.0 req/s), FlexLLM sustains finetuning throughputs of 9.4K, 3.7K, and 3.2K tokens/s for the same models, translating to a 2.5$\times$-6.8$\times$ improvement over the 75% vLLM - 25% LlamaFactory setup.
8.2 GPU Scheduling
This section compares FlexLLM’s co-serving mechanism with two commonly used GPU scheduling strategies: temporal sharing and spatial sharing. Temporal sharing interleaves the execution of inference and finetuning tasks over time, as illustrated in Figure 1. Spatial sharing simultaneously launches inference and finetuning computations using separate CUDA resources. Both strategies are implemented on top of FlexLLM by replacing the co-serving mechanism with their respective scheduling strategies. To ensure a fair comparison, all memory optimizations are enabled in each baseline.
Figure 11 shows the results. For temporal sharing, interleaving one inference iteration with one finetuning iteration would violate the SLOs for nearly all inference requests. This is because each inference iteration is expected to complete within tens of milliseconds, whereas a finetuning iteration requires several seconds. To mitigate this issue, in the temporal sharing experiments, we interleave each finetuning iteration with $n$ inference iterations, where $n$ is the inference frequency shown in the figure.
While adopting a low inference frequency in temporal sharing (i.e., frequency = 64) maximizes finetuning throughput, it adversely impacts the SLO attainment and throughput for inference requests. Employing a frequency of 128 enables temporal sharing to match the inference throughput and SLO attainment rate of co-serving. However, it also reduces finetuning throughput by 1.2-1.8$\times$ compared to co-serving.
Temporal sharing allocates separate GPU resources for inference and finetuning, achieving comparable finetuning throughput to co-serving. However, it remains suboptimal in SLO attainment under heavy inference workloads due to the interference between inference and finetuning tasks.
(a) Arrival Inference Requests
(b) System throughput
Figure 12: Case study of FlexLLM’s system throughput for fluctuating inference workload.
8.3 Case Study
This section describes a case study of FlexLLM using a a 10-min interval of the BurstGPT trace with Qwen-2.5-14B model. To ensure compatibility with our experiment environments, we replayed this trace segment and re-scaled its arrival intensity, as previous works have done [89, 29, 8, 54]. In the experiment, each inference request and inference request are sampled from the ShareGPT dataset and the Sky-T1 dataset respectively.
In our case study, as shown in Figure 12, we observed that the arrival rate of the inference requests initially increased to a peak level after around 90 seconds, and then gradually decreased with some peaks. FlexLLM could automatically detect the fluctuations in the workload and improve the ratio of inference tokens (vs fine-tuning tokens) in each iteration’s batch. This significantly increased inference throughput from a few hundreds to 2.25K.
8.4 Memory Optimization
Figure 13: Ablation study of memory optimizations.
FlexLLM introduces a series of memory optimizations to reduce the memory overhead of PEFT finetuning and allow co-serving finetuning with inference workloads. To understand the effectiveness of FlexLLM memory optimizations, we perform an ablation study that incrementally turns off these optimizations (graph pruning, rematerialization, and token-level finetuning) and measures the memory requirements. Figure 13 shows the activation memory requirements of FlexLLM for different finetuning methods on a 70B LLM and a sequence length of 1024. FlexLLM saves $85\%$-$87\%$ memory requirements for activations compared to existing approaches. The major improvement comes from graph pruning. With graph pruning only, FlexLLM achieves $71\%$-$74\%$ activation memory overhead reduction. Rematerialization and token-level finetuning further reduce memory overhead by 00-$8\%$ and $4\%$-$10\%$, respectively.
9 Related Work
ML serving and inference systems.
Numerous ML serving systems [29, 30, 40, 62, 18] have been proposed in recent years to address various challenges such as latency prediction, system scalability, model performance, swapping [10], preemption [16], workload estimation [89], cost-effectiveness [87], and more. These systems include Nexus [64], which focuses on batching and scheduling multiple small DNN models within a GPU for higher execution throughput, and Gpulet [15], which improves resource assignment and utilization by creating multiple virtual GPUs within a single physical GPU. However, these systems are mainly designed for small CNN models. Recently, LLM serving systems attracts lots of attention [52], such as vLLM [39], Sarathi [7], and SpecInfer [53]. Among them, AlpaServe [45] extends the temporal sharing approach to LLMs via model parallelism across GPUs. Some concurrent PEFT serving systems like Punica [13] and S-LoRA [66] follow the design of PetS [97] and further improve it by leveraging new GPU kernels for latest PEFT methods like LoRA [35]. However, most of these existing works ignore the resource utilization issue in LLM inference systems. To the best of our knowledge, FlexLLM is the first PEFT co-serving system that concurrently executes finetuning and inference tasks and improves resource utilization without affecting inference latency.
GPU resource sharing for ML.
GPU resource sharing is an emerging field of study focusing on the efficient allocation and management of GPU resources among multiple ML users or applications. Lyra [42] allocates dedicated resources for training and serving with dynamic adjustment. Gandiva [81] and Antman [82] take a time-slicing approach to switch between different training jobs for better cluster utilization. GSLICE [21] targets the over- and under-provisioning problems in multi-steam and MPS approaches by developing an adaptive GPU resource allocation and batching scheme for tasks with different SLO constraints. Although these approaches can not directly solve the resource utilization problem in PEFT service, we believe GPU resource sharing will be increasingly important as GPUs become more powerful.
10 Conclusion
This paper proposes FlexLLM, a system for co-serving parameter-efficient finetuning and inference of large language models. We observe the distinct GPU resource utilization features between finetuning and inference workloads and propose a holistic PEFT serving system that supports to co-serve finetuning requests without affecting inference latency.